# Heeji Yoon1*  Jaewoo Jung1*  Junwan Kim2*Hyungyu Choi3 Heeseong Shin1Sangbeom Lim4Honggyu An1Chaehyun Kim1Jisang Han1 Donghyun Kim4

### v1

@@ 5.7 Qualitative Results
Figure 5: Qualitative comparison of baseline and VIRAL. The left part presents PCA visualizations of intermediate representations, demonstrating that VIRAL yields more structured, semantically meaningful visual embeddings. The right part illustrates instance counting and spatial relation tasks, highlighting scenarios where VIRAL correctly answers questions while the baseline fails.
@@
Furthermore, by aligning internal visual representations with robust vision foundation models (VFMs), the semantic quality of intermediate representations is significantly enhanced. This improvement is clearly evidenced in the PCA visualizations shown in Figure 5. We apply PCA to the visual representations obtained from the 16-th layer of Ours and LLaVA-1.5-7B, where our method yields more structured and semantically coherent embeddings compared to the baseline. These visualizations highlight that our alignment strategy effectively guides the model to preserve critical visual details, thereby facilitating better fine-grained visual comprehension. Additional visualizations are provided in Appendix B.1 and B.2.
@@ 6 Conclusion
In this work, we propose VIRAL, a simple yet effective regularization strategy that aligns the internal visual representations of MLLMs with those from pre-trained vision foundation models. Our approach helps preserve fine-grained visual semantics that are often discarded under text-only supervision, thereby enabling more accurate spatial reasoning and object grounding. Extensive experiments across diverse benchmarks validate the effectiveness and generality of our method, showing that visual representation alignment improves both performance and training efficiency in multimodal learning.
@@ Appendix A Additional implementation details
Visual foundation models.
We use a diverse set of pretrained VFMs to supervise internal visual representations. DINOv2 (Oquab et al., 2023), CLIP (Radford et al., 2021), and Depth Anything v2 (Yang et al., 2024) (DAv2) are used as patch size 14 models, while RADIO-v2.5 (Heinrich et al., 2025) and SAM (Kirillov et al., 2023) are used as patch size 16 models. To match the 576 visual tokens produced by CLIP-ViT-L/14 at $\times336$ resolution in LLaVA-1.5 (Liu et al., 2024), we adopt the same resolution for patch size 14 models and resize inputs to $\times384$ for patch size 16 models. For SAM, which expects $\times1024$ inputs, we pad the interpolated features to $\times1024$ and crop them to the region corresponding to the original image, following AM-RADIO (Ranzinger et al., 2024) to avoid quality degradation.
@@
The cosine similarity $\mathrm{sim}(\mathbf{x},\mathbf{y})$, as done in previous works (Hong & Kim, 2021; Hong et al., 2024c; 2022b; 2022a; b; a; Cho et al., 2021; 2022; An et al., 2025), is computed as following $\mathrm{sim}(\mathbf{x},\mathbf{y})=\frac{\mathbf{x}^{\top}\mathbf{y}}{\lVert\mathbf{x}\rVert_{2}\lVert\mathbf{y}\rVert_{2}}$. To balance the alignment loss $\mathcal{L}_{\mathrm{VRA}}$ with the language modeling loss $\mathcal{L}_{\mathrm{LM}}$, we set $\lambda=0.5$ by default.
@@
Figure A1: Visualization of patch random permutation experiments.
@@
For our patch permutation experiment, we adopt the analysis pipeline originally proposed in (Qi et al., 2025). Specifically, we begin by extracting image features $z$ from the vision encoder using $z=V_{\psi}(I)$, where $I$ is the input image. Here, $z\in\mathbb{R}^{N\times H}$, with $N$ denoting the number of visual tokens and $H$ the dimensionality of the vision encoder features. Before processing the vision features $z$ with the vision-language projector $P_{\phi}(\cdot)$ and language model ${LM}_{\theta}(\cdot)$, we apply a random permutation on the order of the visual tokens $N$, which is shown in the visualization of Figure A1. This makes it extremely difficult to understand the visual attributes of the image, enabling us to evaluate how much the MLLM was understanding and utilizing the visual attributes originally available in the image.
Experiments with different vision encoders.
Our method builds on the strong visual representations provided by vision foundation models (VFMs), raising a natural question: does the performance gain stem mainly from these enhanced representations rather than our proposed regularization strategy? To explore this, we conduct a controlled experiment using the same LLaVA-7b-1.5 architecture, substituting its original CLIP vision encoder with SigLIPv2 (Tschannen et al., 2025). We use the google/siglip2-large-patch16-384 variant, which has a comparable parameter count to the original CLIP model. Results show that while stronger vision encoders do improve overall performance, alignment with rich visual features still degrades when supervision is provided only via text. This misalignment is effectively mitigated by our proposed $\mathcal{L}_{\mathrm{VRA,\ell}}$ loss, leading to a similar performance boost.
Appendix B Additional Visualizations
B.1 Layer-wise Internal Representations
@@
We present PCA visualizations of the intermediate visual representations from all layers of LLaVA-1.5-7B and VIRAL in Figure A2, enabling a layer-wise comparison of their representational structures. A qualitative comparison with the baseline reveals that visual representation alignment regularizes the MLLM’s internal visual features, leading to more semantically coherent and structured representation, especially in the middle and later layers where meaningful vision understanding emerges.
B.2 Visual Representations with Different VFMs
@@
In addition to Figure 5, we qualitatively present in Figure A3 PCA visualizations of how internal visual representations evolve when aligned with different VFMs. Compared to the baseline representation from LLaVA-1.5-7B, VFM features exhibit more semantically structured organization. Aligning the MLLM’s internal representations with these VFM features distills such structure, enabling the model to refer to enhanced and more coherent visual representations.
B.3 Attention map visualizations.
@@
In Figure A4, we provide visualizations of text-to-image cross-attention maps in the MLLM to qualitatively support the attention analysis from the main paper. Compared to the baseline, the model trained with our method exhibits improved attention behavior by focusing more accurately and locally on regions relevant to the given multimodal context. This observation aligns well with the spatial entropy analysis in Figure 4, where models trained with visual representation alignment show more focused and discriminative attention patterns.
@@
Figure A2: Layer-wise PCA visualizations of visual representations from (a) LLaVA-1.5-7B and (b) ours.
@@
Figure A3: PCA visualizations of 16-th layer visual representations aligned with different VFMs: CLIP, DINOv2, SAM, DAv2, and RADIO.
Figure A4: Cross-attention map comparison for vision centric tasks.

### v2

@@ 5.6 Qualitative Results
Figure 5: Qualitative comparison of baseline and VIRAL. The first column shows the input image–question pairs, and the next two present LLaVA-1.5 and VIRAL results with PCA visualizations and answers. VIRAL yields structured embeddings and correct answers on counting and spatial tasks where the baseline fails.
@@
Furthermore, by aligning internal visual representations with robust vision foundation models (VFMs), the semantic quality of intermediate representations is significantly enhanced. This improvement is clearly evidenced in the PCA visualizations shown in Figure 5. We apply PCA to the visual representations obtained from the 16-th layer of Ours and LLaVA-1.5-7B, where our method yields more structured and semantically coherent embeddings compared to the baseline. These visualizations highlight that our alignment strategy effectively guides the model to preserve critical visual details, thereby facilitating better fine-grained visual comprehension. Additional visualizations are provided in Appendix F.1 and F.2.
5.7 Training Efficiency
Figure 6: Training Efficiency. Performance with $\mathcal{L}_{\mathrm{VRA}}$ (solid) evaluated every 1K steps, averaging accuracies on CV-Bench2D and MMVP. Models trained with $\mathcal{L}_{\mathrm{VRA}}$ achieve faster convergence. Dashed lines represent converged performance of baseline.
To further showcase the additional benefits of VIRAL, we evaluated vision-centric benchmarks, including CV-Bench2D and MMVP, and averaged the accuracy at every 1K training step from the total of 5.2K training steps of the visual instruction tuning stage in Figure 6. Across three CLIP-based models (CLIP-Vicuna-1.5-7B, CLIP-Qwen2.5-7B, CLIP-Vicuna-1.5-13B), the models trained with VIRALshow that convergence is faster and quickly surpasses the baseline performance in 3K steps. Since our method introduces only about a 3% overhead in total training time, these earlier performance gains may translate into improved scalability.
@@ 6 Conclusion
In this work, we propose VIRAL, a simple yet effective regularization strategy that aligns the internal visual representations of MLLMs with those from pre-trained vision foundation models. Our approach helps preserve fine-grained visual semantics often discarded under text-only supervision, thereby enabling more accurate spatial reasoning and object grounding.
@@ Appendix A Additional implementation details
All experiments in this paper are conducted on four NVIDIA A100 GPUs (40 GB each).
Vision foundation models.
We use a diverse set of pretrained VFMs to supervise internal visual representations. DINOv2 (Oquab et al., 2023), CLIP (Radford et al., 2021), and Depth Anything v2 (Yang et al., 2024) (DAv2) are used as patch size 14 models, while RADIO-v2.5 (Heinrich et al., 2025) and SAM (Kirillov et al., 2023) are used as patch size 16 models. To match the 576 visual tokens produced by CLIP-ViT-L/14 at $\times336$ resolution in LLaVA-1.5 (Liu et al., 2024a), we adopt the same resolution for patch size 14 models and resize inputs to $\times384$ for patch size 16 models. For SAM, which expects $\times1024$ inputs, we pad the interpolated features to $\times1024$ and crop them to the region corresponding to the original image, following AM-RADIO (Ranzinger et al., 2024) to avoid quality degradation.
@@
The cosine similarity $\mathrm{sim}(\mathbf{x},\mathbf{y})$, as done in previous works, is computed as following $\mathrm{sim}(\mathbf{x},\mathbf{y})=\frac{\mathbf{x}^{\top}\mathbf{y}}{\lVert\mathbf{x}\rVert_{2}\lVert\mathbf{y}\rVert_{2}}$. To balance the alignment loss $\mathcal{L}_{\mathrm{VRA}}$ with the language modeling loss $\mathcal{L}_{\mathrm{LM}}$, we set $\lambda=0.5$ by default.
Benchmark settings.
To demonstrate the effectiveness of VIRAL, we evaluate it on widely used benchmarks including CV-Bench, MMVP, What’s Up, POPE, MME, and MM-Star. We only use the 2D subset of CV-Bench, as 3D tasks are beyond our scope. For simplicity, we report overall accuracy on CV-Bench2D instead of separately averaging $\mathrm{ADE20K}$ and $\mathrm{COCO}$. For MMVP, we follow its standard evaluation protocol using pair accuracy, but for stability, we report the average accuracy over 10 runs. For POPE, we evaluate on $\mathrm{COCO}$ following LLaVA and report the average accuracy across the “random” and “popular” subsets. For What’s Up, we report the average accuracy between $\mathrm{COCO_{\text{one}}}$ and $\mathrm{COCO_{\text{two}}}$. For MME, we report MMEEN along with the sum of the perception and cognition categories, and for MM-Star, we follow their standard evaluation protocols.
@@
Figure 7: Visualization of patch random permutation experiments.
@@
For our patch permutation experiment, we adopt the analysis pipeline originally proposed in (Qi et al., 2025). Specifically, we begin by extracting image features $z$ from the vision encoder using $z=V_{\psi}(I)$, where $I$ is the input image. Here, $z\in\mathbb{R}^{N\times H}$, with $N$ denoting the number of visual tokens and $H$ the dimensionality of the vision encoder features. Before processing the vision features $z$ with the vision-language projector $P_{\phi}(\cdot)$ and language model ${LM}_{\theta}(\cdot)$, we apply a random permutation on the order of the visual tokens $N$, which is shown in the visualization of Figure 7. This makes it extremely difficult to understand the visual attributes of the image, enabling us to evaluate how much the MLLM was understanding and utilizing the visual attributes originally available in the image.
Appendix B Extended exploration of the pilot study
Figure 8: Extended exploration of the pilot study.
In Section 4, we demonstrated that MLLMs exhibit progressive visual information loss across layers, and that preserving such information can enhance their visual understanding (Figure 8-(b,c)). In this section, we compare two additional strategies for preserving visual information: a residual connection with the raw encoder feature prior to projection (pre-projection) as a direct approach to feature re-injection (Figure 8-(d)), and our proposed visual representation alignment with the projected features (post-projection) provided to the language model (Figure 8-(e)).
Table 4: Benchmark performance of the pilot study.
POPE CV-Bench2D MMVP
Baseline 85.70% 56.82% 28.20%
(b) 87.17% 57.51% 26.67%
(c) 88.10% 59.53% 29.33%
(d) 85.47% 53.62% 19.33%
(e) 86.99% 57.23% 28.53%
Residual connection with pre-projection features.
Our investigation leverages pre-projection features—raw encoder features $\mathbf{z}$ prior to the projector—through a direct residual connection to mitigate visual information loss within the language model, where a lightweight adapter $P_{\phi’}(\cdot)$ is employed for dimensional compatibility. As illustrated in Figure 8-(d), we conduct one such experiment that re-injects $\mathbf{z}_{i}$ into $\mathbf{e}^{\mathrm{img}}_{\ell,i}$ such that
$\mathbf{e}^{\mathrm{img}}_{\ell,i}\leftarrow\mathbf{e}^{\mathrm{img}}_{\ell,i}+P_{\phi^{\prime}}(\mathbf{z}_{i}).$ (7)
However, as shown in Table 4-(d), this approach generally performs worse than the baseline. This is because the raw encoder features, which have not passed through the pre-trained projector, are not sufficiently aligned with language features (Liu et al., 2023), and their direct residual connection consequently disrupts vision–language alignment in the intermediate layers. These findings suggest that incorporating external features into the internal visual pathway of LLMs requires more careful design.
Visual representation alignment with post-projection features.
Next we further explore aligning the intermediate visual representation with the post-projection features, as shown in Figure 8-(e). Here, we follow the same experimental setting as in Section 4.3, while $\mathcal{L}_{\mathrm{VRA}}$ is defined as:
$\mathcal{L}_{\mathrm{VRA}}=-\,\frac{1}{N}\sum_{i=1}^{N}\mathrm{sim}\!\left(P_{\pi}(\mathbf{e}^{\mathrm{img}}_{\ell,i}),\,P_{\phi}(\mathbf{z}_{i})\right).$ (8)
The results presented in Table 4-(e) indicate that this approach generally improves performance over the baseline on vision-centric benchmarks, yet underperforms compared to leveraging raw features from the vision encoder. This may be attributed to the insufficient preservation of visual information in the post-projection features compared to the raw encoder outputs (Verma et al., 2024; Cha et al., 2024; Li et al., 2025b).
Appendix C Additional ablation studies
Table 5: Ablation study on key design components.
VFM Ladyer Index Objective CV-Bench2D MMVP What’s Up POPE MME
Baseline 56.82% 28.20% 40.13% 85.70% 1650.21
Ablation studies on different multi-layer targets
DINOv2 16 Cos. Sim. 59.67% 33.33% 48.55% 88.32% 1694.52
DINOv2 151715-17 Cos. Sim. 59.32% 28.00% 47.17% 87.61% 1639.72
DINOv2 141814-18 Cos. Sim. 49.62% 22.55% 42.58% 87.90% 1444.32
Ablation studies on different alignment objectives
DINOv2 16 Cos. Sim. 59.67% 33.33% 48.55% 88.32% 1694.52
DINOv2 16 Relation 58.83% 26.60% 49.05% 87.58% 1674.30
Number of target layers.
To investigate the effective number of target layers, we evaluate multi-layer targets around the 16th—specifically $\pm 1$ (15–17) and $\pm 2$ (14–18) ranges—and observe that applying alignment solely at the 16th layer achieves the best performance. These findings highlight that aligning visual representations at a specific pathway responsible for visual representation processing, rather than uniformly across multiple layers, is more effective in enhancing the visual understanding capabilities of MLLMs. Based on this observation, we adopt the 16th layer as the default alignment target with DINOv2.
Alignment objectives.
We investigate the impact of different feature alignment objectives during instruction tuning. Specifically, we compare the performance of models trained with a feature relation alignment objective, as a substitute for the proposed direct visual representation alignment loss. Here, the alignment objective is defined as a mean squared error (MSE) loss between the self-similarity matrices of the VFM features and the transformed intermediate representations, which effectively distills the structural relationships among visual features following recent approaches (Zhang et al., 2025a; Bolya et al., 2025). As shown in Table 5, we find that simple cosine similarity-based alignment loss yields higher performance, and adopt it as our default strategy for alignment.
Appendix D Comparison with other training objectives
Table 6: Comparison with reconstructive objective.
Language Vision Objective CV-Bench2D MMVP What’s Up POPE MMStar MME
Model Encoder
Vicuna-1.5-7B CLIP Baseline 56.82% 28.20% 40.13% 85.70% 33.93% 1650.21
ROSS (Default) 54.24% 29.73% 43.57% 88.19% 34.73% 1648.87
ROSS (Middle) 56.05% 31.40% 45.98% 88.21% 33.53% 1647.27
VIRAL 59.67% 33.33% 48.55% 88.32% 33.93% 1694.52
We compare our method with ROSS (Wang et al., 2024), which applies a reconstructive objective to the final hidden state of the visual representations. To isolate the sources of improvement, we implement two variants under identical experimental conditions: ROSS (Default), reproducing the original method, and ROSS (Middle), which applies the same objective to an intermediate layer (16th layer as in our configuration for target supervision).
Table 6 reveals several key findings that validate our approach. First, the critical importance of intermediate layer supervision—a contribution of our work—is evidenced by ROSS (Middle) outperforming ROSS (Default), particularly on vision-centric benchmarks. Although both ROSS (Default) and ROSS (Middle) show improvements over the baseline which also shows the importance of providing supervision to the visual pathways, the superiority of ROSS (Middle) over ROSS (Defaults) confirms our hypothesis that supervising visual information flow at strategically chosen intermediate layers, rather than naively at the model’s output, yields superior performance gains.
Second, and more fundamentally, our method significantly outperforms both ROSS variants across all benchmarks. This performance gap stems from a crucial distinction in objectives: while ROSS employs reconstruction-based objectives that excel at preserving low-level fidelity, such approaches are inherently less suited for capturing the higher-level semantic abstractions required by complex reasoning tasks (Zhang et al., 2023; Tong et al., 2024a). In contrast, our direct alignment with pretrained vision foundation models provides richer semantic supervision that better bridges the vision-language gap.
These results demonstrate that our method’s superiority arises from two synergistic contributions: (1) the strategic placement of supervision at critical intermediate layers, and (2) the use of semantically-rich alignment signals from vision foundation models rather than reconstruction-based objectives. Together, these design choices enable more effective visual representation learning for multimodal understanding.
Appendix E Applicability of VIRAL to other MLLM architectures
Figure 9: Visual alignment of LLaVA-NeXT (Liu et al., 2024b). Layer-wise alignment between visual tokens in MLLM and vision encoder features, measured by CKNNA and averaged across representations from tiled image splits.
Recent MLLMs employ various strategies to handle inputs with dynamic resolutions, including dynamically adjusting the sequence length of visual tokens (Bai et al., 2023b) or dividing high-resolution images into independently encoded tile grids (Liu et al., 2024b; Chen et al., 2024c). The latter approach preserves the original image resolution and is commonly adopted to capture fine-grained visual details.
To investigate whether VIRAL can be applied to such recent MLLM paradigms, we examine if our core motivation—mitigating vision information loss—remains relevant within this tiled image processing strategy. Figure 9 demonstrates a decline in alignment scores between input visual features and layer-wise visual representations across the layers in LLaVA-NeXT (Liu et al., 2024b), as measured using CKNNA (Huh et al., 2024). This shows similar patterns observed in Figure 2(d), suggesting that VIRAL can be applied orthogonally to such techniques and has the potential to similarly enhance fine-grained visual understanding in MLLMs designed for dynamic resolution handling.
Appendix F Additional visualizations and results
F.1 Layer-wise internal representations
@@
We present PCA visualizations of the intermediate visual representations from all layers of LLaVA-1.5-7B and VIRAL in Figure 10, enabling a layer-wise comparison of their representational structures. A qualitative comparison with the baseline reveals that visual representation alignment regularizes the MLLM’s internal visual features, leading to more semantically coherent and structured representation, especially in the middle and later layers where meaningful vision understanding emerges.
We present PCA visualizations of the intermediate visual representations from all layers of LLaVA-1.5-7B and VIRAL in Figure 10, enabling a layer-wise comparison of their representational structures. A qualitative comparison with the baseline reveals that visual representation alignment regularizes the MLLM’s internal visual features, leading to more semantically coherent and structured representation, especially in the middle and later layers where meaningful vision understanding emerges.
F.2 Visual representations with different VFMs
@@
In addition to Figure 5, we qualitatively present in Figure 11 PCA visualizations of how internal visual representations evolve when aligned with different VFMs. Compared to the baseline representation from LLaVA-1.5-7B, VFM features exhibit more semantically structured organization. Aligning the MLLM’s internal representations with these VFM features distills such structure, enabling the model to refer to enhanced and more coherent visual representations.
F.3 Attention map visualizations
@@
In Figure 12, we provide visualizations of text-to-image cross-attention maps in the MLLM to qualitatively support the attention analysis from the main paper. Compared to the baseline, the model trained with our method exhibits improved attention behavior by focusing more accurately and locally on regions relevant to the given multimodal context. This observation aligns well with the spatial entropy analysis in Figure 4, where models trained with visual representation alignment show more focused and discriminative attention patterns.
Appendix G Limitations
While our method demonstrates general improvements across vision encoders, model scales, and language backbones (Table 1), several considerations remain. First, since VFMs generally produce representations aligned with the spatial grid of the original image, our alignment relies on projection modules that preserve this structure (e.g., linear projection layers, the de facto choice in current MLLMs (Li et al., 2024; Chen et al., 2024d; Bai et al., 2023b; Lu et al., 2024)). Architectures such as Resampler (Alayrac et al., 2022; Wang et al., 2023) or Q-Former (Dai et al., 2023; Li et al., 2023a) disrupt this grid, making our approach less directly applicable. In the same vein, effective alignment also requires that the resulting grid be resolution-adjustable so that the number of visual tokens matches those expected by the MLLM. Also, because our alignment strengthens the semantic utility of each vision token and their relationships, approaches that rely on token pruning to exploit redundancy (Vasu et al., 2025; Wen et al., 2025) may yield reduced benefits when combined with our method. Finally, while our experiments—as well as prior studies—indicate that the middle layers of MLLMs are primarily responsible for fine-grained information, this behavior may not hold universally as more diverse architectures continue to emerge.
Appendix H Use of Large Language Models
In accordance with the ICLR 2026 submission policy, we disclose that Large Language Models were used to assist in grammar correction and polishing of the writing in this paper.
@@ Appendix H Use of Large Language Models
Figure 10: Layer-wise PCA visualizations of visual representations from (a) LLaVA-1.5-7B and (b) VIRAL (Ours).
@@
Figure 11: PCA visualizations of 16th layer visual representations aligned with different VFMs: CLIP (Radford et al., 2021), DINOv2 (Oquab et al., 2023), SAM (Kirillov et al., 2023), DAv2 (Yang et al., 2024), and RADIO (Heinrich et al., 2025).
Figure 12: Cross-attention map comparison for vision centric tasks.